# MoA: Mixture of Sparse Attention for Automatic Large Language Model Compression

### v2

@@ A.1 Main Setup
Baselines. In the setup for our experiment, we adhere to specific configurations outlined in the respective papers. In the case of In StreamingLLM (Xiao et al., 2024c), the initial four tokens remain unmasked, serving as the attention sink, except for the 70b model in Table 4 and the super long setting in Figure 5, where we use 64 tokens as the attention sink. For InfLLM (Xiao et al., 2024a), we adhere to the original configuration by maintaining the same local window size and selected memory size, using 128 initial tokens as specified in their setup. For H2O (Zhang et al., 2023), we ensure the same number of heavy hitter tokens and recent tokens. Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction. StreamingLLM, InfLLM and MoA use sparse prefill.
@@
Models and Benchmarks. Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k (Chiang et al., 2023) can only take in 16k context length, we use the 16k split of LV-Eval benchmark (Yuan et al., 2024), truncating the input to 15500 for model input in Table 4. For the LongBench benchmark (Bai et al., 2023), we use the LongBench-E split, which features a balanced number of data items at every length level. The LongBench dataset is segmented into ranges of 0-4k, 4-8k, and 8k+ tokens. We test each split using the input length truncation thresholds of 3,500, 7,500, and 15,500 tokens, respectively.
@@
Perplexity Evaluation. We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper (Dasigi et al., 2021), MultiNew  (Fabbri et al., 2019), TREC (Li & Roth, 2002; Hovy et al., 2001) and LCC (Mohler et al., 2016), representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM. Following LongBench, the data items are organized as question-answer pairs. The questions and answers are written by humans and come with the dataset. The perplexity is calculated solely on the answer part of the data, demonstrating the model’s coherence in responding to user requests.
@@
Validation Dataset. The validation dataset is used to select the optimal compression plan among the Pareto front solutions during the optimization step. The validation dataset is similarly constructed as the perplexity test dataset, but on the respective validation split of the datasets. $50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper (Dasigi et al., 2021), MultiNew  (Fabbri et al., 2019), TREC (Li & Roth, 2002; Hovy et al., 2001) and LCC (Mohler et al., 2016). The additional $50$ data items from the LongEval (Li et al., 2023) dataset are also added to validate the retrieval ability. For the datasets that do not contain the validation split, namely TREC, MultiNews and LCC, we sample from the test split and ensure different data items with the perplexity evaluation dataset.
@@
MoA Settings. MoA uses the block sparse attention pattern with a block size of 64, where each grid depicted in Figure 3(a) represents a block. The first block of tokens is not masked as the attention sink. For the profile stage, we use the MultiNews (Fabbri et al., 2019) calibration dataset with model response as supervision, as described in Section 5. We use $50\times 3$ data items at 2k, 4k, 8k lengths. The data items are padded to their corresponding length level in order to ensure a unified shape of attention influence tensors for each length level. We adopt block granularity during the profiling stage, calculating the average attention influence within each block to represent the block’s overall influence. The optimization is done with the multi-objective optimization at the same set of lengths. We limit the number of distinct rules to at most two per model layer to ensure inference-time efficiency. Among the Pareto front solutions, we select the one with the lowest perplexity on the validation dataset of length 12k.
A.2 Efficiency Experiment Setup
@@
For MoA, The implementation is based on Huggingface Transformers. During the prefill stage, we use the sparse CUDA kernel designed by us with block size 64. During the decode stage, we modify the KV-Cache implementation to support our heterogeneous elastic rules. Thanks to our fixed sliding-window span during the decode stage, we simply replace the old KV-Cache that exceeds the span with the latest KV-Cache. Our custom decoding CUDA kernel then handles KV-Cache with varying lengths across different attention heads during the decoding process.
For H2O, we use its official efficient implementation, which is based on Flexgen (Sheng et al., 2023). Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction, which requires the attention matrix to be explicitly calculated. It makes H2O’s prefill stage currently incompatible with kernel optimizations like FlashAttention. Therefore, H2O is easy to get OOM (Out-Of-Memory) with large prefill length and increased batch size.
In our efficiency tests across all frameworks, we implemented a simple optimization at the language modeling head (lm head) during the prefill stage. Specifically, after the final layer of the transformers, we compute the logits—these are the raw outputs that are transformed into probabilities—for only the last token. This selective computation avoids generating these probabilities for preceding tokens, substantially reducing both computational overhead and memory usage. We also set the environment variable PYTORCH_CUDA_ALLOC_CONF to be expandable_segments:True for Hugginface and MoA to mitigate memory fragmentation, allowing larger inference batch size.
Following the performance experiments, we use Vicuna-7B and Vicuna-13B for efficiency tests whenever possible. However, the official efficient implementation of H2O based on Flexgen only supports OPT (Zhang et al., 2022). Therefore, we use OPT-6.7b and OPT-13b models for H2O in Table 9 for comparison.
A.3 Ablation Study Setup
@@
In the calibration dataset experiments in Table 2, we intentionally exclude the influence of the validation dataset. We avoid using the validation dataset by profile and optimize solely at 8k length, reducing the multi-objective optimization problem to a single-objective one with only one optimal compression plan instead of a set of Pareto fronts.
Appendix B Additional Experiment Results
B.1 Performance
B.1.1 Overall Performance
@@
B.1.2 Long-Context Retrieval
@@
We conduct a detailed experiment to test the retrieval ability of different attention methods across various attention spans and input lengths.
@@
Following previous work (Chen et al., 2023; Tworkowski et al., 2023), we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold. As shown in Figure 7(a), StreamingLLM and H2O achieve effective context lengths of no more than 2k tokens beyond their attention spans. In contrast, MoA expands its effective context length to approximately 3.9×3.9$\times$3.9 × its attention span before reaching up to the 12k limit of the original model. Figure 7(b) further shows that at a fixed input length of 8k, MoA reaches over 0.9 retrieval accuracy with just 25% density, whereas StreamingLLM and H2O require 100% and 75% density, respectively.
B.1.3 Long-Context Understanding
@@
We conduct experiments with various densities on the LV-Eval benchmark (Yuan et al., 2024). As shown in Figure 8, MoA constantly outperforms the uniform static attention baseline StreamingLLM at various densities, demonstrating the effectiveness of our heterogeneous elastic rules.
@@
We detailed the respective scores for LongBench and LV-Eval in Table 7 and Table 8. The number in the bracket of Table 8 indicates the number of sub-datasets for the category.
B.2 Efficiency
B.2.1 Throughput Breakdown
@@
In this section, we explain the decode throughput breakdown for MoA, compared to the baseline comprising Huggingface with FlashAttention2. The observed increase in throughput, as shown in Table 5, primarily stems from three aspects:
@@
B.2.2 Automatic Compression Pipeline Overhead
Table 10: Compression overhead for various stages of MoA across models with differing parameter sizes, reported as the amount of GPU $\times$ latency, except when only one GPU is used. Larger models necessitate more GPUs due to model parallelism. All stages utilize GPUs, except for the Optimize stage, which uses the CPU.
@@ B.2.2 Automatic Compression Pipeline Overhead
Table 11: Progressive compression overhead for various stages of MoA with respect to different parameter sizes and calibration (validation) dataset sizes.
@@
On the Intel(R) Xeon(R) Platinum 8358 2.60 GHz CPU, the optimization concludes within approximately 25 minutes. Typically, this phase generates around 10 compression plans. Validating each one of the compression plans takes about 4 minutes, totaling around 40 minutes.
We also show the progressive compression overhead for MoA in Table 11.
B.3 Ablation Study
B.3.1 Calibration Dataset
Table 12: Performance comparison on various test sets, using different calibration sets. Tested on Vicuna-7B model. The result is tested with 50% density MoA on LongBench (Bai et al., 2023) 0-4k split.

### v3

@@ B.1 Main Setup
Baselines. In our experimental setup, we adhere to the specific configurations outlined in the respective papers. In the case of StreamingLLM (xiao2023streamingLLM), the initial four tokens remain unmasked, serving as the attention sink, except for the 70b model in Table 4 and the super long setting in Figure 5, where we use 64 tokens as the attention sink. For InfLLM (xiao2024infllm), we adhere to the original configuration by maintaining the same local window size and selected memory size, using 128 initial tokens as specified in their setup. For H2O (Zhang2023H2O), we ensure the same number of heavy hitter tokens and recent tokens. Note that H2O uses original dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction. StreamingLLM, InfLLM and MoA use efficient prefill. Notably, in MoA and StreamingLLM, the KV-Cache length equals the attention span during the efficient prefill stage. In contrast, H2O uses the original dense prefill. Besides, H2O and InfLLM require additional computations to dynamically determine the KV-Cache.
@@
Models and Benchmarks. For long-context retrieval, we use LongEval (lmsys2023longeval) to test key-value retrieval accuracy with 100 data items per length level. For long-context understanding, we use LV-Eval (yuan2024lveval) and LongBench (bai2023longbench), which include 11 and 13 sub-datasets, respectively. Since vicuna-7b-v1.5-16k and vicuna-13b-v1.5-16k (vicuna2023) can only take in 16k context length, we use the 16k split of LV-Eval benchmark (yuan2024lveval), truncating the input to 15500 for model input in Table 4. For the LongBench benchmark (bai2023longbench), we use the LongBench-E split, which features a balanced number of data items at every length level. The LongBench dataset is segmented into ranges of 0-4k, 4-8k, and 8k+ tokens. We test each split using the input length truncation thresholds of 3,500, 7,500, and 15,500 tokens, respectively. Efficiency experiments measure the decode throughput on a single A100-80GB GPU at maximum batch sizes of respective methods.
@@
Perplexity Evaluation. We construct a comprehensive yet concise test set by sampling $50\times 4$ data items for each length level from the test split of four long-context understanding datasets: Qasper (Dasigi2021Qasper), MultiNews (Fabbri2019MultiNews), TREC (li-roth-2002-trec1; hovy-etal-2001-trec2) and LCC (lcc), representing the question answering, summarization, few-shot learning, and code completion abilities of the LLM. Following LongBench, the data items are organized as question-answer pairs. The questions and answers are written by humans and come with the dataset. The perplexity is calculated solely on the answer part of the data, demonstrating the model’s coherence in responding to user requests.
@@
Validation Dataset. The validation dataset is used to select the MoA configuration among the Pareto front solutions during the optimization step. The validation dataset is similarly constructed as the perplexity test dataset, but on the respective validation split of the datasets. $50\times 4$ data items are sampled from the same four long-context understanding datasets: Qasper (Dasigi2021Qasper), MultiNews (Fabbri2019MultiNews), TREC (li-roth-2002-trec1; hovy-etal-2001-trec2) and LCC (lcc). The additional $50$ data items from the LongEval (lmsys2023longeval) dataset are also added to validate the retrieval ability. For the datasets that do not contain the validation split, namely TREC, MultiNews and LCC, we sample from the test split and ensure different data items with the perplexity evaluation dataset.
@@
MoA Settings. MoA uses the block sliding-window attention pattern with a block size of 64, where each grid depicted in Figure 3(a) represents a block. The first block of tokens is not masked and serves as the attention sink. For the profiling stage, we use the MultiNews (Fabbri2019MultiNews) calibration dataset with the model response as supervision, as described in Section 5. We use $50\times 3$ data items at 2k, 4k, and 8k lengths. The data items are padded to their corresponding length level to ensure a unified shape of attention-influence tensors for each level. We adopt block granularity during profiling, calculating the average attention influence within each block to represent the block’s overall influence. For hyperparameter search space $\alpha$ and $\beta$, we use 6 values for $\alpha$ and 9 values for $\beta$, creating a search space of 54 pairs for each attention head. $\alpha$ is uniformly sampled from the range [2048,8192][-2048,8192], and $\beta$ is uniformly sampled from $[0,1]$. The resulting attention span lengths are clipped to the range between 0 and the current input length. The optimization is done with the multi-objective optimization at the same set of lengths. We limit the number of distinct rules to at most two per model layer to ensure inference-time efficiency. Among the Pareto front solutions, we select the one with the lowest perplexity on the validation dataset of length 12k.
B.2 Efficiency Experiment Setup
@@
For MoA, the implementation is based on HuggingFace Transformers. During the prefill stage, we use the sliding-window attention CUDA kernel that we designed with a block size of 64. During the decode stage, we modify the KV-Cache implementation to support our heterogeneous elastic rules. Thanks to our fixed sliding-window span during decoding, we simply replace the old KV-Cache that exceeds the span with the latest KV-Cache. Our custom decoding CUDA kernel then handles KV-Cache with varying lengths across different attention heads during decoding.
For H2O, we use its official efficient implementation, which is based on FlexGen (sheng2023flexgen). Note that H2O uses dense prefill since it relies on the column sum of the attention matrix to calculate the importance of every token for KV-Cache eviction, which requires the attention matrix to be explicitly calculated. This makes H2O’s prefill stage currently incompatible with kernel optimizations like FlashAttention. Therefore, H2O is prone to OOM (Out-Of-Memory) with large prefill lengths and increased batch sizes.
In our efficiency tests across all frameworks, we implemented a simple optimization at the language modeling head (lm head) during the prefill stage. Specifically, after the final layer of the transformers, we compute the logits—these are the raw outputs that are transformed into probabilities—for only the last token. This selective computation avoids generating these probabilities for preceding tokens, substantially reducing both computational overhead and memory usage. We also set the environment variable PYTORCH_CUDA_ALLOC_CONF to expandable_segments:True for HuggingFace and MoA to mitigate memory fragmentation, allowing larger inference batch sizes.
Following the performance experiments, we use Vicuna-7B and Vicuna-13B for efficiency tests whenever possible. However, the official efficient implementation of H2O based on Flexgen only supports OPT (zhang2205opt). Therefore, we use OPT-6.7b and OPT-13b models for H2O in Table 11 for comparison.
B.3 Ablation Study Setup
@@
In the calibration dataset experiments in Table 2, we intentionally exclude the influence of the validation dataset. We avoid using the validation dataset by profile and optimize solely at 8k length, reducing the multi-objective optimization problem to a single-objective one with only one optimal MoA configuration instead of a set of Pareto fronts.
B.4 Input Format And Examples
We list the prompt format and input examples used in our primary experiments and datasets. Dashed lines are included only for illustration clarity and are not part of the texts given to the LLMs.
Format 1. LongEval Below is a record of lines I want you to remember. Each line begins with ’line <line index>’ and contains a ’<REGISTER_CONTENT>’ at the end of the line as a numerical value. For each line index, memorize its corresponding <REGISTER_CONTENT>. At the end of the record, I will ask you to retrieve the corresponding <REGISTER_CONTENT> of a certain line index. Now the record start: line delightful-incandescence: REGISTER_CONTENT is <19147> line cloistered-presence: REGISTER_CONTENT is <8862> Now the record is over. Tell me what is the <REGISTER_CONTENT> in line cloistered-presence? I need the number.
Format B.4 illustrates the input format for the LongEval (lmsys2023longeval) retrieval benchmark. The instruction indicating which line to retrieve is provided after a lengthy context containing massive lines of register contents to remember.
Format 2. Needle-In-A-Haystack (NIAH) People who are powerful but uncharismatic will tend to be disliked. Their power makes them a target for criticism that they don’t have the charisma to disarm. That was Hillary Clinton’s problem. The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day. It also tends to be a problem for any CEO who is more of a builder than a schmoozer. What is the best thing to do in San Francisco?
Format B.4 depicts the input format for another common retrieval benchmark, Needle-In-A-Haystack (NIAH) (kamradt2024NIAH). The NIAH test comprises a single "needle" sentence that commonly does not fit into an irrelevant context. The model tries to answer the question based on this needle sentence.
Format 3. MultiNews Calibration Dataset You are given several news passages. Write a one-page summary of all news. <News1> <News2> Now, write a one-page summary of all the news. Summarization
Format B.4 demonstrates the input format for our calibration dataset. The long-context MultiNews dataset (Fabbri2019MultiNews) consists of multiple news documents. The context includes a prompt instructing the original dense model to generate a summarization for these news articles, reflecting long-range dependencies and model alignment. The generated summarization serves as supervision during the cross-entropy loss calculation at the profiling stage.
Appendix C Additional Experiment Results
C.1 Performance
C.1.1 Overall Performance
@@
C.1.2 Long-Context Retrieval
@@
LongEval Retrieval. We conduct a detailed experiment to test the retrieval ability of different attention methods across various attention spans and input lengths with the LongEval (lmsys2023longeval) dataset.
@@
Following previous work (Chen2023PI; Tworkowski2023FocusedTrans), we quantify effective context length as the maximum input length where retrieval accuracy remains above a 90% threshold. As shown in Figure 7(a), StreamingLLM and H2O achieve effective context lengths of no more than 2k tokens beyond their attention spans. In contrast, MoA expands its effective context length to approximately $3.9\times$ its attention span before reaching up to the 12k limit of the original model. Figure 7(b) further shows that at a fixed input length of 8k, MoA reaches over 0.9 retrieval accuracy with just 25% density, whereas StreamingLLM and H2O require 100% and 75% density, respectively.
Figure 8: The Needle-In-A-Haystack (NIAH) retrieval accuracy using different attention methods across 8k to 256k input lengths on Llama-3-8B model. All efficient attention methods employ a 50% density.
Needle-In-A-Haystack (NIAH) Retrieval. We also conduct the retrieval task using the Needle-In-A-Haystack (NIAH) dataset (kamradt2024NIAH). As shown in Figure 8, MoA achieves perfect retrieval accuracy across input lengths ranging from 8k to 256k. In comparison, StreamingLLM demonstrates a limited effective context length, while InfLLM exhibits reduced retrieval accuracy within 64k input lengths. Notably, H2O and InfLLM are unable to complete tests at extreme lengths due to Out-Of-Memory and Out-Of-Time errors. These findings align with the results observed in the LongEval benchmark throughout the paper.
C.1.3 Long-Context Understanding
@@
We conduct experiments with various densities on the LV-Eval benchmark (yuan2024lveval). As shown in Figure 9, MoA constantly outperforms the uniform static attention baseline StreamingLLM at various densities, demonstrating the effectiveness of our heterogeneous elastic rules.
@@
We detailed the respective scores for LongBench and LV-Eval in Table 7 and Table 8. The number in the bracket of Table 8 indicates the number of sub-datasets for the category. MoA achieves comprehensive performance comparable to the original dense model, as well as H2O that requires higher efficiency cost. In contrast, StreamingLLM and InfLLM display inconsistent performance: it sometimes surpasses the original model in some tasks, while suffering noticeable degradation in others.
C.1.4 Longer-Context Generalization
Retrieve Acc. $\uparrow$
Attention 32k 64k 128k 256k
SnapKV 1.00 0.88 0.71 0.33
PyramidKV 1.00 0.85 0.62 0.37
MoA 1.00 0.92 0.83 0.46
Table 9: Retrieval accuracy at longer lengths for more recent baselines, tested at 50% density.
We compare the retrieval accuracy with more recent works SnapKV (li2024snapkv) and PyramidKV (cai2024pyramidkv) on context lengths of 32K to 256K. As shown in Table 9, MoA constantly outperforms the two latest baselines at longer contexts.
C.1.5 Instruction-following Generation
Attention Length-controlled Win Rate $\uparrow$ Standard Error
Original 8.84 0.53
H2O 9.66 0.55
InfLLM 5.76 0.42
StreamingLLM 7.96 0.49
MoA 9.83 0.57
Table 10: Length-controlled win rate and its standard error of Vicuna-7B with different attention mechanisms on AlpacaEval 2.0 benchmark. All efficient attention methods employ 50% density during decoding.
We evaluate MoA ’s performance on general instruction-following tasks using the AlpacaEval 2.0 benchmark (alpacaEvalRepo; dubois2024alphacaEvalLength). Following the official setup, we compare the model’s output with gpt4_turbo using the standard weighted_alpaca_eval_gpt4_turbo evaluator, which leverages the gpt-4-1106-preview model. The benchmark consists of inputs and outputs with average lengths of approximately 50 and 450 tokens, respectively. To accommodate the short input lengths while maintaining a density of around 50% during generation, we set the expected total token length to 512 and adjust hyperparameters across all methods accordingly.
Thanks to its elastic design, MoA employs the same configuration used in experiments with input lengths ranging from 4k to 256k. As shown in Table 10, MoA achieves the highest length-controlled win rate, outperforming both efficient attention baselines and the original model.
C.2 Efficiency
C.2.1 Memory and Throughput Breakdown
@@
Table 11 highlights the memory efficiency of MoA compared to H2O and FlashAttention2 on 7B and 13B models. Notably, H2O runs into Out-Of-Memory (OOM) issues at longer input lengths. In contrast, MoA achieves a significant reduction in memory consumption, using $1.2$ to $1.4\times$ less memory compared to FlashAttenion2.
We further explain the decode throughput breakdown in Table 5, compared to the baseline comprising Huggingface with FlashAttention2. The observed increase in throughput primarily stems from four aspects:
@@
GPU Kernel Design. We customize MoA GPU kernel using CUDA to support heterogeneous attention patterns with high efficiency.
C.2.2 Efficiency Results for Longer Input
Min. Total Total Throughput
Model Size Framework Attention #GPU Throughput Memory (GB) per GPU
7B vLLM PagedAttention 2 30.2 142.0 15.1
FlexGen H2O >>8 - OOM -
HuggingFace InfLLM 1 6.1 47.7 6.1
HuggingFace StreamingLLM 1 19.8 43.9 19.8
HuggingFace FlashAttention2 2 4.3 85.6 2.2
HuggingFace MoA 1 20.3 44.0 20.3
13B vLLM PagedAttention 2 21.5 142.0 10.8
FlexGen H2O >>8 - OOM -
HuggingFace InfLLM 1 4.3 78.6 4.3
HuggingFace StreamingLLM 1 14.0 64.6 14.0
HuggingFace FlashAttention2 2 3.0 130.6 1.5
HuggingFace MoA 1 14.7 63.4 14.7
Table 12: Runtime efficiency at 128k input length across different methods on Vicuna-7B and 13B models. All efficient attention methods use 50% density. Decode throughput (tokens per second) is measured with a batch size of 1, using the minimum number of A100-80GB GPUs required for testing. H2O encounters OOM error with 8 GPUs.
We evaluate the runtime efficiency of Vicuna-7B and 13B models at a 128k input length with a single batch size. Thanks to the reduced KV-Cache, MoA efficiently processes 128k input using only one A100 GPU, whereas FlashAttention2 and vLLM baselines require at least two GPUs to handle a single request. As shown in Table 12, MoA achieves a $4.7$-$4.9\times$ decode speedup compared to FlashAttention2, while using half the number of GPUs. Additionally, it demonstrates a $1.9$-$2.1\times$ reduction in GPU memory usage. Compared to vLLM, which utilizes tensor parallelism, MoA delivers $1.3$-$1.4\times$ higher throughput per GPU, alongside significant memory savings.
C.2.3 Energy Efficiency Result
[…]
Energy Per Token (J) Power (W)
Framework Attention 4k 8k 16k 4k 8k 16k
Huggingface FlashAttention2 2.98 5.93 12.1 350 354 359
Huggingface MoA 0.34 0.62 1.21 330 322 315
Table 13: Per token energy consumption and average GPU power with different attention mechanisms at various sequence lengths, using Vicuna-7B on a single A100 GPU.
C.2.4 Automatic Search Pipeline Overhead
@@
Table 14: Search overhead for various stages of MoA across models with differing parameter sizes, reported as the amount of GPU $\times$ latency, except when only one GPU is used. Larger models necessitate more GPUs due to model parallelism. All stages utilize GPUs, except for the Optimize stage, which uses the CPU.
@@
Table 15: Overheads for various stages of MoA with respect to different parameter sizes and calibration (validation) dataset sizes.
@@
On the Intel(R) Xeon(R) Platinum 8358 2.60 GHz CPU, the optimization concludes within approximately 25 minutes. Typically, this phase generates around 10 MoA configurations. Validating each one of the configuration takes about 4 minutes, totaling around 40 minutes.
We also show the overhead of each stages in Table 15.
C.3 Ablation Study
C.3.1 Calibration Dataset